A homeowner takes a 30-year fixed-rate mortgage for $145,000 at 8.05 percent. After twelve years, the homeowner sells the house and pays off the remaining principal. How much is the principal payment? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1
we have to use financial calculator to solve this problem
First we have to compute 12 year EMI payment
Put in calculator
FV 0
PV 145000
I 8.05%
N 30
compute PMT ($12,940.79)
therefore annual payment = $12,940.79
we can solve now using Amortization table to know the principal value at the 15th year.
i ii=i*8.05% iii iv=i+ii-iii
year Beginning balance Interest EMI Closing balance
1 145000 11672.5 $12,940.79 $143,731.71
2 $143,731.71 11570.403 $12,940.79 $142,361.33
3 $142,361.33 11460.087 $12,940.79 $140,880.63
4 $140,880.63 11340.891 $12,940.79 $139,280.73
5 $139,280.73 11212.099 $12,940.79 $137,552.04
6 $137,552.04 11072.94 $12,940.79 $135,684.20
7 $135,684.20 10922.578 $12,940.79 $133,665.99
8 $133,665.99 10760.112 $12,940.79 $131,485.31
9 $131,485.31 10584.568 $12,940.79 $129,129.09
10 $129,129.09 10394.892 $12,940.79 $126,583.20
11 $126,583.20 10189.947 $12,940.79 $123,832.36
12 $123,832.36 9968.5049 $12,940.79 $120,860.08
13 $120,860.08 9729.2361 $12,940.79 $117,648.53
14 $117,648.53 9470.7063 $12,940.79 $114,178.44
15 $114,178.44 9191.3648 $12,940.79 $110,429.02
Therefore closing balance = $110,429.02
so principal to be paid = $110,429.02